OpenStudy (anonymous):

Question about completing the square? This is my problem: 6m^2 + 5m + -3 = 0 I have no clue where to start. I know how to start a quadratic equation, but completing a square has really confused me.

OpenStudy (anonymous):

Hello?

OpenStudy (whpalmer4):

First, divide through by 6 so that the coefficient of the m^2 term is 1. \[x^2+5x-3=0\] \[x^2 + \frac{5}{6}x - \frac{1}{2}= 0\] Now we move the last term to the right by adding 1/2 to each side, and then to complete the square, we take half of the coefficient of the middle term, square it, and add that to both sides: \[x^2+\frac{5}{6}x + \frac{1}{2}+(\frac{5}{6*2})^2=\frac{1}{2} +(\frac{5}{6*2})^2\] Having done that, we can write the left side as a square (we just completed it!) \[(x+\frac{5}{12})^2 = \frac{1}{2}+\frac{25}{144}\] and now the rest is just taking the square root of both sides: \[x+\frac{5}{12} = \sqrt{\frac{72}{144}+\frac{25}{144}}\] \[x+\frac{5}{12}=\pm\frac{\sqrt{79}}{12}\] \[x=\frac{-5\pm\sqrt{79}}{12}\]

OpenStudy (anonymous):

Thanks so much for your help! I really appreciate this.